solve for x: 5+(3x+1)=5(x+1)+2
Answer by checkley79(3341)

5+(3x+1)=5(x+1)+2
5+3x+1=5x+5+2
3x-5x=5+2-5-1
-2x=1
x=1/-2
x=-1/2 ans.
Proof:
5+(3*-1/2+1)=5(-1/2+1)+2
5-1.5+1=5*1/2+2
6-1.5=2.5+2
4.5=4.5